OpenStudy (anonymous):

Solve by factoring: x2+4x+3=0

mathslover (mathslover):

\[\large{x^2+4x+3=0}\] \[\large{x^2+x+3x+3=0}\] right?

OpenStudy (anonymous):

yes

mathslover (mathslover):

now: \[\large{x(x+1)+3(x+1)=0}\] \[\large{(x+3)(x+1)=0}\] right?

mathslover (mathslover):

now either x+3 = 0 or x + 1 = 0 can you solve these 2 eqn now?

OpenStudy (anonymous):

yes, i just subtract 3 and get x=-3 and subtract 1 and get x=-1 , right?

mathslover (mathslover):

right very right ..

mathslover (mathslover):

good ..

OpenStudy (anonymous):

so the answer is x= -3 &-1

mathslover (mathslover):

yes